Find the missing value. the rate is 17%, the time is 4 years, the simple interest is $10,268.00, i have to find the principal.
1
Expert's answer
2012-11-29T08:07:55-0500
from the problem we have equation
x*1.7^4 = 10.268
where x is principal
so x=10.268/1.7^4 = 1 229.39